# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v1

@@ Abstract
Information retrieval in hour-long videos presents a significant challenge, even for state-of-the-art Vision-Language Models (VLMs), particularly when the desired information is localized within a small subset of frames. Long video data presents challenges for VLMs due to context window limitations and the difficulty of pinpointing frames containing the answer. Our novel video agent, FALCONEye [^1] Webpage: https://cplou99.github.io/FALCONEye/, combines a VLM and a Large Language Model (LLM) to search relevant information along the video, and locate the frames with the answer. FALCONEye novelty relies on 1) the proposed meta-architecture, which is better suited to tackle hour-long videos compared to short video approaches in the state-of-the-art; 2) a new efficient exploration algorithm to locate the information using short clips, captions and answer confidence; and 3) our state-of-the-art VLMs calibration analysis for the answer confidence. Our agent is built over a small-size VLM and a medium-size LLM being accessible to run on standard computational resources. We also release FALCON-Bench, a benchmark to evaluate long (average $>1$ hour) Video Answer Search challenges, highlighting the need for open-ended question evaluation. Our experiments show FALCONEye’s superior performance than the state-of-the-art in FALCON-Bench, and similar or better performance in related benchmarks.
Figure 1: Overview of our meta-architecture FALCONEye designed to Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs. A VLM pre-process the video generating captions from small video clips. A LLM iteratively refines the search by focusing on the most promising clips using the captions. The exploration follows a hierarchical reasoning: initially, a broad set of low-resolution frames is analyzed by the VLM for each promising clip, progressively narrowing down to fewer frames with higher resolution as the search concentrates on smaller temporal clips. The LLM uses captions, question semantics, answer completion, and confidence scores to determine whether to continue exploring or end the exploration. Once a high-confidence answer is found, or the search reaches a predefined threshold, FALCONEye outputs the final answer, its confidence score, and the corresponding temporal interval.
@@ 1 Introduction
Vision-Language Models (VLMs) have shown remarkable performance in answering complex, fine-grained questions about images and short videos [17, 4, 32], but solving similar tasks on long videos remains an open problem [5, 53]. Besides the intrinsic problem complexity, processing and encoding all video frames detailed information is too costly and impractical [54]. Our work emphasis is on analyzing very long videos that contain the answer on a tiny set of frames, a needle in a haystack problem, that we will designate as Video Answer Search (VAS) task. Thus, the challenge is to locate the specific content necessary to answer the question in an efficient and practical way.
Large-Language Models (LLMs), and consequently VLMs, are constrained by their context window size, which limits the number of tokens they can process in a single pass. This constraint is not critical to process a 300-page book or an image. However, when dealing with videos, only a small subset of frames (typically in the order of dozens or hundreds) can be sampled and encoded to keep the total number of tokens within the context window size [52, 47]. This sampling strategy may work for short videos, but is likely to fail for extremely long videos, where using only a few frames is likely to miss a significant part of relevant information. Therefore, research is increasingly focused on adapting VLMs to work with long-form videos [37, 38, 6].
First, Video Question Answering (VQA) benchmarks are extending video durations and question complexity [5, 3, 42]. However, the focus is mainly on multiple-choice questions, where the LLM’s ability to eliminate options may be even more crucial than its ability to infer the answer from the sampled frames  [35]. In this work, we release FALCON-Bench, specifically designed to evaluate VAS in (90 publicly available) long-form videos, focusing on (more than 580) open-ended questions that require to pinpoint specific information, measuring both accuracy and localization.
Second, several VLMs specifically designed for long videos have emerged [38, 37, 4], leveraging smart frame sampling and token compression techniques to maximize the amount of sampled visual information that fits within its context window. However, they often lack the precision required for VAS task when only a tiny fraction of the video contains the answer, as they provide a global representation instead of bootstrapping resources in the most relevant video segments. In this context, we introduce FALCONEye, a video agent to tackle VAS task. Through an iterative approach that integrates a VLM and a LLM, FALCONEye leverages captions, question-answers semantics, and confidence scores to optimize localization, avoiding irrelevant clips and concentrating resources on exploring temporally relevant segments until finding the answer (Figure 1).
@@
Overall, our work presents the following contributions: 1) FALCON-Bench Benchmark: The first benchmark specifically designed for the VAS task. Unlike other VQA benchmarks, it focuses on open-ended questions that require pinpointing answers within tiny temporal regions of one-hour-long videos. 2) Meta-architecture for VAS (FALCONEye): A system that integrates VLM and LLM capabilities to emulate human-like VAS behavior. It leverages a novel search algorithm that reasons to iteratively focus on video clips most likely to contain the answer. 3) Evaluation of state-of-the-art VLMs calibration: A study assessing the calibration of leading VLMs in VQA tasks. Both benchmark and model have been built under the lmms-eval framework [51] in order to ease the reproducibility of the results and future evaluations with other models and benchmarks. FALCONEye and FALCON-Bench represent a significant advancement in VAS and, VQA in general, setting a new standard for precision-guided exploration in long-form video answer search task.
@@ 2 Related work
Table 1: Comparison of our FALCON-Bench and other popular benchmarks for VQA task in videos. We focus on labeling, number (#) of videos and question (QAs), its average and total duration in seconds (s) and hours (h), respectively, and the theme of the videos. Lastly, we analyze if the evaluation is oriented to multiple-choice (MC) or open-question (O), if the benchmark contains questions focused on single detail (SD) aspects of the video, and, if the ground truth (GT) answer contains the GT time interval in which the answer is contained.
Benchmark Labels #Videos #QAs Avg. Dur. (s) Total Dur. (h) Theme Category Eval. SD GT TIME
ActivityNet-QA [49] Human 5,800 58,000 180 40 Everyday Life O
NeXT-QA [46] Human 5,440 52,044 44 12.22 Everyday Life MC/O
MVBench [21] Auto 4,000 4,000 16 17.78 Life, Human Action, Movie MC
EgoSchema [28] Auto 5,031 5,031 180 251.55 Life, Human Action, Egocentric MC
MovieChat-1K [38] Human 130 1,950 500 18.06 Movie MC/O
HourVideo [5] Human 500 12,976 2,742 380 Life, Human Action, Egocentric MC
Event-Bench [14] Auto+Human 1294 2,190 101 36.46 Life, Egocentric, Fun, Cartoon MC (21%)
LONGVIDEOBENCH [44] Human 3,763 6,678 473 494.49 Life, Movie, Knowledge, News MC (30%)
MLVU [53] Human 1,334 2,593 720 267.8 Movie, Surveillance, Egocentric MC/O (64%)
Video-MME [15] Human 900 2,700 1,017 254.47 TV, Sports, Artistic, Life MC (15%)
LVBench [42] Human 103 1,600 4,101 117.3 Sports, TV shows, animations MC (14%)
InfiniBench [3] Auto 1219 108,200 3,155 1,068.3 Movie, TV shows MC/O (14%)
FALCON-Bench (Ours) Human 92 576 4,738 121.1 Soccer, Movie, Egocentric MC/O (100%)
@@
VLMs emerged to extend LLMs’ capabilities to vision, with the seminal work of LLaVA [25], which integrated a pre-trained CLIP visual encoder, ViT-L/14 [33], to extract visual tokens, with an LLM (Vicuna [8]) for multimodal text and image understanding. This integration appears in other open-source architectures which improved the alignment and processing of visual features, such as BLIP-2 [20], LLaVA-1.5 [24], mPLUG-Owl2 [47], and LLaMa3.2-Multimodal [1]. More recently, Qwen2-VL [41] introduced a VLM with a ViT capable of processing images at arbitrary resolutions, dynamically converting them into a variable number of visual tokens. Notably, many companies have integrated increasingly powerful VLMs into their proprietary long-context chatbot systems, including OpenAI’s GPT-4V [32], Google DeepMind’s Gemini [11] or Anthropic’s Claude [2]. VLMs excel at image questions, but extending these capabilities to videos remains an open challenge due to the large number of tokens required to represent video content. VLM strategies for video extends VLMs to process multiple frames as input [22, 19, 21, 17, 52] employing a uniform video sampling strategy to extract and process up to 64 video frames, to keep the amount of visual tokens within the LLM’s context window. More specific long-video solutions target more efficient and sophisticated video representations. MovieChat [38] combines short- and long-term token-based memories to build a global video representation, LongVU [37] presents a spatiotemporal adaptive compression to retain the most relevant parts for question answering, and LongVILA [6] introduces Multi-Modal Sequence Parallelism to extend the number of video frames processed from 8 to 2048. Apollo [54] conducts an extensive study on strategies to enhance long-video understanding. Recently, Qwen2.5-VL [4] extends its dynamic resolution capabilities to the temporal dimension, allowing processing up to 768 frames. Unlike these methods focused on achieving a global understanding of the entire video, our approach exploration-based strategy bootstraps resources in the most relevant video segments for the question.
VQA Benchmarks.
@@
Existing Video Question Answering (VQA) benchmarks evaluate VLMs capabilities for a wide range of video comprehension tasks. Early benchmarks focus on short videos and questions to understand temporal actions [46], human actions [49], egocentric videos [28] or movies [21]. Similar to ours, Event-Bench [14] focuses on single-detail events occurred within short temporal windows but uses one-minute-long videos. Recently, several VQA benchmarks [5, 44, 53, 15] have extended the temporal duration of videos to around 10 minutes. Unlike these benchmarks, which primarily focus on Multiple-Choice Questions (MCQs), MovieChat-1K [38] is centered on open-ended questions (OQs). LVBench [42] and InfiniBench [3] further push this challenge with, up to one hour, longer videos. Some of these later benchmarks include a percentage of single-detailed questions, where the answer is found within a short time window of the video, but only LONGVIDEOBENCH [44] and InfiniBench evaluate a small fraction of these questions in open-ended mode. Our FALCON-Bench is designed to address the challenges of single-detail questions and open-ended responses in very long videos. This is the first benchmark to provide ground truth time intervals that contain the answers, ensuring that models can be evaluated to find and localize answers. Table 1 presents a comparison of all these benchmarks.
Meta-architectures over VLMs for VQA.
Many strategies have emerged to tackle the most challenging VQA scenarios, often defined as meta-architectures agnostic to specific models [45]. Long Story Short [10] leverages a LLM that summarizes the video through clip captions, selects the most relevant clips, generates potential answers from their captions, and combines the likelihood of the option letter tokens with the similarity between question and frame embeddings to determine the final response. VideoAgent [43], which shares some similarities with our approach, introduces a three-step algorithm within a training-free meta-architecture. It integrates a VLM to generate clip captions and an LLM to produce answers with confidence scores, selecting the most relevant frames based on the similarity between CLIP embeddings of the frames and the question features. MoReVQA [29] and ProViQ [9] proposes a decomposed multi-stage, modular reasoning framework based on ViperGPT [39], a library of API modules to assemble executable programs for video reasoning. Ranasinghe et al. [34] present the MVU framework, which injects object-centric information—extracted via object detection and tracking models—into an LLM. TimeChat [36] fine-tunes a VLM built on LLaMa-2 by introducing two novel components: a timestamp-aware frame encoder and a sliding video Q-Former, enhancing fine-grained temporal understanding. SeViLA [48] follows a sequential approach, chaining a pre-trained image-language model (BLIP-2) to first localize relevant frames and then answer questions based on them. As most approaches are designed for short videos, longer VQA benchmarks [5] often adopt the Socratic method [50], where an LLM answers questions using captions pre-extracted by a VLM from short clips.

### v2

@@ Abstract
Finding information in hour-long videos is a challenging task even for top-performing Vision Language Models (VLMs), as encoding visual content quickly exceeds available context windows. To tackle this challenge, we present FALCONEye [^1] Webpage: https://cplou99.github.io/FALCONEye/, a novel video agent based on a training-free, model-agnostic meta-architecture composed of a VLM and a Large Language Model (LLM). FALCONEye answers open-ended questions using an exploration-based search algorithm guided by calibrated confidence from the VLM’s answers. We also introduce the FALCON-Bench benchmark, extending Question Answering problem to Video Answer Search—requiring models to return both the answer and its supporting temporal window for open-ended questions in hour‑long videos. With just a 7B VLM and a lightweight LLM, FALCONEye outscores all open‑source 7B VLMs and comparable agents in FALCON-Bench. It further demonstrates its generalization capability in MLVU benchmark with shorter videos and different tasks, surpassing GPT‑4o on single‑detail tasks while slashing inference cost by roughly an order of magnitude.
Figure 1: Overview of our meta-architecture FALCONEye designed to Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs. A VLM pre-processes the video generating captions from small video clips. A LLM iteratively refines the search by focusing on the most promising clips using the captions. The exploration follows a hierarchical reasoning: initially, a broad set of low-resolution frames is analyzed by the VLM for each promising clip, progressively narrowing down to fewer frames with higher resolution as the search concentrates on smaller temporal clips. The LLM uses captions, question semantics, answer completion, and confidence scores to determine whether to continue exploring or end the exploration. Once a high-confidence answer is found, or the search reaches a predefined threshold, FALCONEye outputs the final answer, its confidence score, and the corresponding temporal interval.
@@ 1 Introduction
Vision Language Models (VLMs) have demonstrated remarkable performance on complex, fine-grained questions over images and short videos [llava-onevision, Qwen2.5-VL, openai2023gpt4v]. However, extending these capabilities to long-form videos remains an open challenge [chandrasegaran2024hourvideo, zhou2024mlvu, xiao2025videoqa]. Our work emphasis is on analyzing very long videos and single-detail questions whose answers are contained on a tiny set of frames, a needle in a haystack problem. Not only that, we go a step further by pushing the task boundaries—focusing on open-ended questions and requiring models to return both the answer and visual evidence of the answer in the form of a temporal window where the answer is found. This forces the agent to perform a deeper reasoning, limiting the number of hallucinations or random guesses deduced purely from the question or options without visual confirmation [rana2025mvu]. We refer to this extended setting as Video Answer Search (VAS) task.
Unlike general video understanding—where a global, embedded representation of the entire video may suffice for coarse-level tasks—VAS demands precise information retrieval within a large temporal space. Current VLMs typically address long video inputs by uniformly sampling frames (typically in the order of dozens or hundreds) until reaching the context window size limit [zhang2024llavanext-video, ye2024mplug]. While this strategy may support high-level summarization or scene classification, it often fails when the answer lies in a small, temporally-localized segment [xiao2025videoqa]. In VAS, locating a single relevant moment among thousands of frames requires not just compression, but active reasoning and targeted exploration—capabilities that current VLMs still lack.
Recent efforts have focused on fine-tuning VLMs for long-form video understanding [park2024too, wang2025videotree, shen2024longvu, song2024moviechat, longvila, Qwen2.5-VL, shu2025video, shu2025video]. These models aim to encode more frames by optimizing sampling and token compression strategies to fit within the LLM’s context window. While this enables a more complete global representation of the video, it comes at the cost of temporal resolution and precision—making them insufficient for tasks like VAS, where identifying a short, specific segment is crucial. To bridge this gap, several LLM-based agents such as such as VideoAgent [fan2024videoagent] or VideoTree [wang2025videotree] use meta-architectures combining multiple models to extract key frames—often via CLIP features—and answer multiple-choice questions.
In contrast, we introduce FALCONEye, a novel video agent based on a training-free, model-agnostic meta-architecture composed solely of a VLM and an LLM. It is capable of answering single-detail, open-ended questions in hour-long videos by following a video search algorithm guided by calibrated confidence from the VLM’s outputs—avoiding irrelevant clips and focus computational resources on exploring the most promising clips. Figure 1 illustrates FALCONEye’s dynamic and iterative approach.
Overall, our work presents the following contributions: 1) FALCON-Bench: The first benchmark specifically designed for VAS task. It focuses on open-ended (OE) questions that require locating answers in tiny temporal regions of one-hour-long videos, filling the challenging gap left by existing VQA benchmarks regarding accuracy and temporal answer localization for OE questions in longer videos.
@@
2) Novel video agent (FALCONEye): A system to emulate human-like behavior to answer single-detail questions in videos, integrating VLM and LLM capabilities in a training-free and model agnostic meta-architecture. It leverages our novel search algorithm, that reasons to iteratively focus on video clips most likely to contain the answer. FALCONEye and FALCON-Bench have been built under the lmms-eval framework [zhang2024lmmseval] to facilitate reproducibility and future evaluations. They are a significant advancement in VAS and, VQA in general, setting a new standard for precision-guided exploration in long-form video.
3) Evaluation of state-of-the-art VLMs calibration: A study assessing the calibration of leading VLMs in VQA.
@@ 2 Related work
VLMs emerged to extend LLMs’ capabilities to vision, with the seminal work of LLaVA [liu2023llava], which integrated a pre-trained CLIP visual encoder (ViT-L/14 [radford2021learning]) to extract visual tokens, with an LLM (Vicuna [chiang2023vicuna]) for multimodal text and image understanding. This integration appears in other open-source architectures which improved the alignment and processing of visual features, such as BLIP-2 [li2023blip], LLaVA-1.5 [liu2023improvedllava], mPLUG-Owl2 [ye2024mplug], and LLaMa3.2-Multimodal [llama3_2].
More recently, Qwen2-VL [Qwen2-VL] introduced a VLM with a ViT capable of processing images at arbitrary resolutions, dynamically converting them into a variable number of visual tokens. Many companies have integrated increasingly powerful VLMs into their proprietary long-context chatbot systems, including OpenAI’s GPT-4V [openai2023gpt4v], Google DeepMind’s Gemini [google2023gemini] or Anthropic’s Claude [anthropic2024claude]. VLMs excel at image questions, but extending these capabilities to videos remains an open challenge due to the large number of tokens required to represent video content.
@@
VLM strategies for video extend VLMs to process multiple frames as input [lin2023video, llava-next, li2024mvbench, llava-onevision, zhang2024llavanext-video] employing a uniform video sampling strategy to extract and process up to 64 frames, to keep the visual token amount within the LLM’s context window. Specific long-video solutions target efficient and sophisticated video representations. MovieChat [song2024moviechat] builds a global video representation combining short- and long-term token-based memories, Video-XL-Pro [liu2025video] and LongVU [shen2024longvu] present a spatiotemporal adaptive compression to retain the most relevant parts for question answering, and LongVILA [longvila] introduces Multi-Modal Sequence Parallelism to extend the number of video frames processed from 8 to 2048. Apollo [zohar2024apollo] conducts an extensive study on strategies to enhance long-video understanding. Differently, ReKV [di2025rekv] is a training-free mechanism that retrieves relevant information from a video key-value cache to enable efficient streaming VQA. Recently, Qwen2.5-VL [Qwen2.5-VL] extends its dynamic resolution capabilities to the temporal dimension, allowing processing up to 768 frames. Unlike these methods focused on answering by means of global complete video understanding, we focus on finding a short clip that contains the answer and returning the answer.
Table 1: FALCON-Bench compared to popular video VQA benchmarks: number of videos (#V) and questions (#QAs); average video duration in seconds (s); multiple-choice (MC) or open-question (O) in the evaluation (underlined if one of them is dominant in the benchmark); inclusion of questions focused on single detail (SD) video aspects; provided GT time interval ($\Delta t$) where the answer is contained; VAS task evaluation. A key distinction is wether QA and ground truth (QA-GT) are human annotations or automatically extracted from available video description (Auto).
Benchmark QA-GT #V #QAs Dur.(s) Eval. SD $\Delta t$ VAS
ActivityNet-QA [yu2019activitynet] Human 5,800 58,000 180 O
MVBench [li2024mvbench] Auto 4,000 4,000 16 MC
EgoSchema [mangalam2023egoschema] Auto 5,031 5,031 180 MC
Event-Bench [du2024towards] Both 1,294 2,190 101 MC (21%)
LongVideoBench [wu2024longvideobench] Human 3,763 6,678 473 MC (30%)
LongVideoBench-L [wu2024longvideobench] Human 966 - 1408 MC (30%)
Video-MME [fu2024video] Human 900 2,700 1,017 MC (15%)
VIDEO MME-Long [fu2024video] Human 300 900 2466,7 MC (15%)
HourVideo [chandrasegaran2024hourvideo] Human 500 12,976 2,742 MC
LVBench [wang2024lvbench] Human 103 1,600 4,101 MC (14%)
NeXT-QA [xiao2021next] Human 5,440 52,044 44 MC/O
MovieChat-1K [song2024moviechat] Human 130 1,950 500 MC/O
MovieChat-1K $10^{\prime}$ [song2024moviechat] Human 19 - >600>600 MC/O
CG-Bench [chen2025cgbench] Human 1,219 12,129 1,624 MC/O (100%)
MLVU [zhou2024mlvu] Human 1,730 3,102 930 MC/O (64%)
MLVU $30^{\prime}$ [zhou2024mlvu] Human 125 - >1800>1800 MC/O (64%)
InfiniBench [ataallah2024infinibench] Auto 1,219 108,200 3,155 MC/O (14%)
QAEgo4D [Baermann_2022_qaego4d] Auto 1,325 14,513 495,1 MC/O (100%)
FALCON-Bench Human 92 576 4,738 MC/O (100%)
@@
VQA Benchmarks. Existing Video QA benchmarks evaluate VLMs capabilities for plenty of video comprehension tasks. Early benchmarks focus on short videos and questions to understand temporal actions [xiao2021next], human actions [yu2019activitynet], egocentric videos [mangalam2023egoschema] or movies [li2024mvbench]. Similar to ours, Event-Bench [du2024towards] focuses on single-detail events occurred within short temporal windows but uses one-minute-long videos. Recently, several VQA benchmarks [chandrasegaran2024hourvideo, wu2024longvideobench, zhou2024mlvu, fu2024video] extend the temporal duration of videos to around 10 minutes. Unlike these benchmarks, primarily focused on Multiple-Choice Questions (MCQs), MovieChat-1K [song2024moviechat] is centered on open-ended questions (OQs). LVBench [wang2024lvbench] and InfiniBench [ataallah2024infinibench] further push this challenge with longer videos, up to 1 hour. Some of these later benchmarks include a percentage of single-detailed questions, with the answer located within a short video time window, but only LongVideoBench [wu2024longvideobench] and InfiniBench evaluate a small fraction of the questions in open-ended mode. Recently, CG-Bench [chen2025cgbench] introduced a benchmark focused on single-detail questions and medium-duration videos (average of 26 minutes per video), being the first to provide ground-truth time intervals for answers and to evaluate state-of-the-art VLMs on the accuracy and temporal localization separately. QAEgo4D [Baermann_2022_qaego4d] also introduces a task close to VAS in egocentric videos based on the Ego4D NQL narrations and annotations [grauman2022ego4d]. However, it contains videos of average length 8 minutes and questions automatically generated from the narration, usually focused on the main actor or object being interacted. In contrast, our FALCON-Bench takes the challenge a step further: it moves to one-hour-long videos, focuses on open-ended questions, and evaluates the VAS task—requiring models to jointly return both the answer and its corresponding supporting temporal window. Table 1 highlights FALCON-Bench more challenging (longer and OE focused) benchmark for VAS task, including Human annotations.
VQA Agent–style meta-architectures. Recent works have built agents for video QA combining an LLM with other architectures (e.g., VLMs, feature extractors). On one hand, there are fine-tuned solutions, e.g., TimeChat [ren2024timechat] fine-tunes a VLM with timestamp-aware encoder and sliding video Q-Former; or SeViLA [yu2023self] chains a fine-tuned BLIP-2 to localise frames before answering. On the other hand, there are very successful training-free pipelines. MoReVQA [min2024morevqa] and ProViQ [choudhury2024video] assemble multi-stage programs with ViperGPT [suris2023vipergpt] modules, while MVU [ranasinghe2024understanding] and VideoAgent2 [ranasinghe2024understanding] inject object-centric tracks into the LLM. Long Story Short [chung2023long] uses a VLM to caption clips, a LLM to select the most relevant ones for a given MCQ and visual-text cosine similarity from CLIP [radford2021clip] features to choose. the predicted option. Similarly, LifelongMemory [wang2024lifelongmemory] leverages an LLM to filter pre-extracted captions and answer MCQs from them. VideoAgent [wang2024videoagent] loops over (i) CLIP-based frame selection (ii) VLM captions of key frames, (iii) LLM answers a MCQ from captions with an uncalibrated confidence score until reaching a confident answer. VideoAgent[fan2024videoagent] couples a VLM captioner [zhao2023learning], a ViCLIP video encoder [wang2023internvid], a GPT‑based LLM, an object detector[zhao2024detrs] and an object tracker [zhang2022bytetrack] to tackle VQA. VideoTree [wang2025videotree] builds a hierarchical tree of key frames (pre-extracted with EVA CLIP-8B [sun2024eva]), captions each node with a VLM and queries the LLM for the final answer. The only agent that only requires a VLM and an LLM is the Socratic baseline [zeng2022socratic, zhang2023simple], where an LLM reasons only over VLM-generated captions. In this line, FALCONEye (i) is a training-free and model-agnostic meta-architecture that only requires a VLM and an LLM; (ii) is the first video agent to bring focus on strategies particularly designed for the more challenging open-ended VQA; (iii) its exploration algorithm is guided by a calibrated confidence estimate derived from the VLM’s open answer, not by an ad-hoc integer from the LLM; (iv) is powered by an adaptive search at clip-level: the agent varies clip length, frame count and resolution on-the-fly instead of relying on key-frame extraction ; (v) unlike existing agents, relying on dense captioning containing the answer, it reasons from very light captions to generate the answer directly from the target visual content.